# Multilingual Performance of a Multimodal Artificial Intelligence System on Multisubject Physics Concept Inventories

### v1

@@ Abstract
We investigate the multilingual and multimodal performance of a large language model-based artificial intelligence (AI) system, GPT-4o, on a diverse set of physics concept inventories spanning multiple languages and subject areas. The inventories taken from the PhysPort website cover the classical physics topics of mechanics, electromagnetism, optics, and thermodynamics as well as relativity, quantum mechanics, astronomy, mathematics, and laboratory skills. Unlike previous text-only studies, we uploaded the inventories as images mirroring what a student would see on paper, assessing the system’s multimodal functionality. The AI is prompted in English and autonomously chooses the language of its response – either remaining in the nominal language of the test, switching entirely to English, or mixing languages – revealing adaptive behavior dependent on linguistic complexity and data availability. Our results indicate some variation in performance across subject areas, with laboratory skills standing out as the area of poorest performance. Furthermore, the AI’s performance on questions that require visual interpretation of images is worse than on purely text-based questions. Questions that are difficult for the AI tend to be that way invariably of the inventory language. We also find large variations in performance across languages, with some appearing to benefit substantially from language switching, a phenomenon similar to code-switching of human speakers. Overall, comparing the obtained AI results to the existing literature, we find that the AI system outperforms average undergraduate students post-instruction in all subject areas but laboratory skills.
@@ I.1 Generative artificial intelligence in physics education
The public availability of Large Language Models (LLMs), like those built on the architecture introduced by Vaswani et al. [1], has unlocked new possibilities across various domains, including education [2, 3]. Since the release of ChatGPT in the fall of 2022 [4], LLMs have surged in popularity, with scholars showcasing their remarkable capabilities. Beyond the wave of enthusiasm generated by human-like responses that have been shown to pass the Turing Test [5] with a majority of human test subjects [6], the GPT series demonstrated proficiency in academic fields such as physics [7]. Both the initial version and later iterations, particularly GPT-4 [8], have achieved impressive results, such as passing standardized exams, excelling in introductory courses and even coming close to passing entire degrees [9, 10, 11, 12, 13].
The technology is increasingly being embraced in physics education [14], offering promising applications for both teaching and learning physics. LLMs have proven valuable for teachers, helping in the creation of tailored materials and tasks, student assessments, and personalized feedback [15, 16, 17], as well as for their training [18]. For students, these models may represent ever-available, patient, and knowledgeable resources [19, 20]. However, these opportunities come with significant risks, particularly the potential for users to overly trust AI when assessing its scientific accuracy [21, 22].
Early publicly available generative AI systems have been limited to processing and generating text-based content only. Thus, earlier studies of AI’s performance on physics tasks, such as concept inventories, were restricted to text-based materials or textual descriptions of visual elements [9, 23]. More recently, multimodal systems, which can also input and output auditory and visual data, have broadened the scope of such studies (e.g., [24]).
In this study, we confront GPT-4o [25], a popular model from OpenAI, with screenshots from physics concept inventories, which reflect the test questions as seen by learners, including the accompanying diagrams, sketches, graphs, and illustrations. Earlier studies using GPT-4 and GPT-4o with the FCI [26], TUG-K [27] and BEMA [28] concept inventories demonstrated promising behavior, but also significant challenges, mainly related to the AI system’s limited ability to interpret visual information [29, 30, 31]. We considered the English version of each selected concept inventory and all their available translations.
@@ I.2 Concept inventories in physics
Examinations in physics generally involve symbolic and numerical calculations, whereas concept inventories are typically different, focusing primarily on conceptual understanding. Scores on these different types of assessment do not necessarily correlate [32]. On the surface, this would work in AI’s favor, as AI systems have been notoriously “bad at math,” which hampered their performance on traditional physics exam questions [9]. However, early investigations of AI’s performance on conceptual tasks suggest that these, too, can present several challenges [33, 10].
@@
Arguably, while many concept inventory questions reflect students’ understanding of core ideas [36], they do not necessarily capture evidence of scientific practices or crosscutting concepts [37] (even though this distinction is debatable [38]). It is thus important to emphasize that our study will assess the equivalent of AI’s abilities and skills in physics, but not evaluate if it has the qualities of a physicist.
@@ I.3 The language problem
Despite all the power and possibilities of using LLMs in education, there are also many limitations. OpenAI’s research on the GPT-4’s performance with different languages, for example, shows inconsistent results across languages [39]. LLMs learn patterns, make predictions, and perform tasks by leveraging large datasets, deep learning algorithms, and reinforcement learning from human feedback [40]. However, due to the disparities in the prevalence, quantity, and quality of information available across languages, there exists a disparity in the resources available for LLM training, which can have an impact on the quality of the outcomes. Nicholas and Bhatia highlight that although LLMs are designed to mitigate the issue of underrepresentation of certain languages in learning data, as of 2023, early LLMs were still predominantly trained on English materials. Although this may be changing as governments and companies attempt to strengthen AI capabilities in their countries (e.g., models like DeepSeek [41] or Qwen [42] developed by companies in China, or the Swiss AI Initiative [43]), it is likely that LLMs remain biased toward the needs of major economies with the financial resources to train AI systems. Thus, this can contribute to the transfer of cultural values and assumptions from English, or languages spoken in other major economies, to languages that lack infrastructure for training AI models. The authors stress the necessity to address this issue by companies, researchers, and governments across the globe [44]. Similar issues are also discussed in the analysis conducted by “Cohere for AI” company in their report titled “The AI Language Gap” [45]. Feng et al. also acknowledge the disparity in LLM performance across different languages, resulting in a gap of approximately 20% between high-resource and low-resource languages [46]. Particularly relevant for our study is a recent finding that a Chinese-trained model performed better on the FCI when prompted in Chinese rather than English [47].
@@
Understanding and using physics-specific language is essential for physics literacy and learning. To address both the linguistic disparities of LLMs and the specialized nature of physics language, it is important to examine how LLMs perform across diverse languages in the context of physics education. While current research highlights the general challenges of multilingual use of LLMs, the intersection of language and subject-specific terminology, such as physics, is understudied and lacks a clear understanding of the current situation. This underscores the importance of evaluating LLMs not only in terms of language proficiency but also in their ability to accurately convey domain-specific knowledge.
Human speakers frequently overcome language disparities by engaging in code-switching [48, 49], switching languages mid-dialogue or even mid-sentence. Within physics education, much of the research on code-switching takes place in bilingual classrooms [50, 51], where students have a native language and a potentially different language of instruction; students tend to use their highest proficiency language when trying to understand what is happening conceptually [51].
I.4 Research questions
The overarching research questions are:
@@ I.4 Research questions
How does a state-of-the-art (2024/25) multimodal AI system perform across different physics concepts?
How does its performance compare to published post-instruction average results at the undergraduate level?
How does language influence the performance of the AI system?
How does the presence or necessity of images influence the performance of the AI system?
@@ II Data set
As data set, we used concept tests published in PhysPort [52], which contains a comprehensive collection of research-based inventories in multiple translations [53]. We included tests that had at least a “bronze star” classification assigned on PhysPort, meaning that they had undergone significant research validation, and for which we could obtain an answer key. Figure 1 shows questions from the Force Concept Inventory in Persian and the Heat and Temperature Conceptual Evaluation in Chinese.
Table 1 shows the languages currently found in PhysPort, as well as the ISO 639 [54] shorthands we will be using throughout the paper; the rightmost column shows the number of inventories we were able to evaluate in each language. While the authors are familiar with some of the languages, the quality of the majority of the translations could not be evaluated.
Tables 3 through 6 show the concept tests included in our investigation, along with their available translations and literature references. For some inventories, more than one version was available. The column “%Post” lists some post-instruction inventory scores reported in the literature for undergraduate-level courses; these are collected best-effort and not necessarily representative (typically and much more systematically, inventory gains are reported [55]). We omitted results with very small sample sizes, as well as reported scores at the graduate-student or post-doctoral level. For subsequent comparisons, the scores found for each inventory were averaged.
The rightmost column “Cat.” shows the subject-area identifier under which we classified the test. Explanations for the abbreviations used in this column are provided in Table 2. These labels are derived from the PhysPort classification, however, we divided electricity and magnetism into EM-F (primarily dealing with fields and potentials), and EM-C (primarily focusing on DC and AC circuits). The multi-subject Next Gen Physical Science Diagnostic (NGPSD) was given no classification.
| Abbreviation | Description | Number |
| AR | Arabic | 4 |
| BN | Bengali | 1 |
| CA | Catalan | 1 |
| CS | Czech | 2 |
| DE | German | 5 |
| EL | Greek | 4 |
| EN | English | 53 |
| ES | Spanish | 16 |
| FA | Persian | 2 |
| FI | Finnish | 7 |
| FIL | Filipino | 1 |
| FR | French | 6 |
| HE | Hebrew | 2 |
| HI | Hindi | 1 |
| HR | Croatian | 1 |
| HU | Hungarian | 1 |
| ID | Indonesian | 5 |
| IS | Icelandic | 1 |
| IT | Italian | 2 |
| JA | Japanese | 9 |
| MS | Malay | 3 |
| NL | Dutch | 1 |
| NO | Norwegian | 1 |
| PA | Punjabi | 1 |
| PL | Polish | 1 |
| PT | Portuguese | 7 |
| RU | Russian | 1 |
| SK | Slovak | 1 |
| SR | Serbian | 1 |
| SV | Swedish | 8 |
| TA | Tamil | 1 |
| TH | Thai | 4 |
| TR | Turkish | 3 |
| UK | Ukrainian | 1 |
| ZH | Chinese | 8 |
@@
EN, ES, JA, PT, SV, ZH | EM-F |
@@
CS, EN | MATH |
| CDPA | Concise Data Processing Assessment | Lab skills (uncertainty in measurement, relationship between functions graphs and numbers) | [62] | 2 | 39 [62] | EN, ES | LAB |
@@
DE, EL, EN, ES, FI, SV, ZH | EM-C |
| DS | Density Survey | Mechanics content knowledge (density) | [69] | 1 | 57 [70] | DE, EN | MECH |
@@
EN, FI, ID, SV | MECH |
@@
AR, BN, CA, CS, DE, EL, EN, ES, FA, FI, FIL, FR, HE, HI, HR, HU, IS, IT, JA, MS, NL, NO, PA, PL, PT, RU, SV, SK, TA, TH, TR, ZH | MECH |
| FMCE | Force and Motion Conceptual Evaluation | Mechanics content knowledge (kinematics, forces, energy, graphing) | [80] | v99 | 55 [81] | EN, ES, ID, JA | MECH |
| FORT | Montana State University Formal Reasoning Test | Scientific reasoning (hypothesis testing, correlational reasoning, probability, control of variables, proportional reasoning) | [82] | 1 | 55 [82] | EN | REAS |
| FTGOT | Four-tier Geometrical Optics Test | Waves / Optics content knowledge (plane mirrors, spherical mirrors, lenses); used as two-tier test | [83] | 1 | 18 [83] | EN, TR | OPT |
@@
EN | MECH |
| GECI | Greenhouse Effect Concept Inventory | Astronomy content knowledge (types of greenhouse gases, types of electromagnetic energy, energy equilibrium balance, greenhouse effect mechanisms, global warming vs. greenhouse effect) | [85] | vC | 55 [85] | EN, JA | AST |
| HTCE | Heat and Temperature Conceptual Evaluation | Thermal / Statistical content knowledge (temperature, phase change, heat transfer, thermal properties of materials) | [86] | 1 | 78 [86] | EN, ZH | THERM |
@@
EN: 26
FR: 31 [87] |
EN, FR | MECH |
| LPCA | Light Phenomena Conceptual Assessment | Waves / Optics content knowledge (reflection, refraction, Snells law, wavelength and frequency, light scattering, electromagnetic spectrum, the human eye) | [88] | 1 | 41 [88] | EN | OPT |
@@
DE, EL, EN, ES, FA, FI, FR, IT, JA, MS, PT, TR | MECH |
@@
EN | EM-F |
| MUQ | Measurement Uncertainty Quiz | Lab skills (calculating error from measurements, accuracy and precision, sources of error) | [97] | 1 | — | EN | LAB |
| MWCS | Mechanical Wave Conceptual Survey | Waves / Optics content knowledge (mechanical waves, wave propagation, wave superposition, wave reflection, standing waves) | [98] | 1 | 47 [99] | EN, ES, TH | OPT |
| 2 | EN, ES | ||||||
@@
AR, EN | AST |
| NGPSD | Next Gen Physical Science Diagnostic | Mechanics content knowledge (magnetism, static electricity, energy, forces, waves and sound, light) | [101] | 2 | — | EN | — |
| PIQL | Physics Inventory of Quantitative Literacy | Mathematics scientific reasoning (proportional reasoning, reasoning with signed quantities, co-variational reasoning) | [102] | 2.2 | 55 [102] | EN | REAS |
@@
| QMCA | Quantum Mechanics Concept Assessment | Modern / Quantum Content knowledge (wave functions, measurement, time dependence, probability, infinite square well, 1D tunneling, energy levels, spins) | [103] | 5.5.7 | 54 [103] | EN, PT | QUANT |
| 6.6.2 | EN | ||||||
@@
EN, ES | QUANT |
| QMS | Quantum Mechanics Survey | Modern / Quantum content knowledge (wave functions, measurement, expectation values, Hamiltonian, time dependence, probability, infinite square well, finite square well, harmonic oscillator, 1D tunneling) | [107] | 18 | 38 [107] | EN | QUANT |
@@
EN | EM-F |
| RCI | Relativity Concept Inventory | Modern / Quantum content knowledge | [111] | 1 | 71 [111] | EN | RELA |
@@
EN, FR | MECH |
| RRMCS | Rotational and Rolling Motion Conceptual Survey | Mechanics content knowledge (rotational kinetic energy, torque, rotational kinematics, moment of inertia) | [115] | 1 | 75 [115] | EN | MECH |
| Title | Full Title | Description | Refs. | Vers. | %Perc | Languages | Cat. |
| SGCE | Symmetry and Gauss’s Law Conceptual Evaluation | Electricity / Magnetism content knowledge (symmetry, electric field, electric flux) | [116] | 1 | 49 [116] | EN | EM-F |
@@
EN, ES, JA | AST |
| STPFaSL | Survey of Thermodynamic Processes and First and Second Laws | Thermal / Statistical content knowledge (first law of thermodynamics, second law of thermodynamics, PV diagrams, reversible processes, irreversible processes) | [118] | short | 37 [118] | EN, ID, PT, ZH | THERM |
| long | EN | ||||||
| TCE | Thermal Concept Evaluation | Thermal / Statistical content knowledge (temperature, heat transfer, phase change, thermal properties of materials) | [119] | 1 | 78 [119] | EN, JA, PT, ZH | THERM |

### v3

@@ Abstract
We investigate the multilingual and multimodal performance of a large language model-based artificial intelligence (AI) system, GPT-4o, using a diverse set of physics concept inventories spanning multiple languages and subject categories. The inventories, sourced from the PhysPort website, cover classical physics topics such as mechanics, electromagnetism, optics, and thermodynamics, as well as relativity, quantum mechanics, astronomy, mathematics, and laboratory skills. Unlike previous text-only studies, we uploaded the inventories as images to reflect what a student would see on paper, thereby assessing the system’s multimodal functionality. Our results indicate variation in performance across subjects, with laboratory skills standing out as the weakest. We also observe differences across languages, with English and European languages showing the strongest performance. Notably, the relative difficulty of an inventory item is largely independent of the language of the test. When comparing AI results to existing literature on student performance, we find that the AI system outperforms average post-instruction undergraduate students in all subject categories except laboratory skills. Furthermore, the AI performs worse on items requiring visual interpretation of images than on those that are purely text-based. While our exploratory findings show GPT-4o’s potential usefulness in physics education, they highlight the critical need for instructors to foster students’ ability to critically evaluate AI outputs, adapt curricula thoughtfully in response to AI advancements, and address equity concerns associated with AI integration.
@@ I.1 Generative artificial intelligence in physics education
The public availability of Large Language Models (LLMs), like those built on the architecture introduced by Vaswani et al. [1], has unlocked new possibilities across various domains, including education [2, 3]. Since the release of ChatGPT in the fall of 2022 [4], LLMs have surged in popularity, with scholars showcasing their remarkable capabilities. Beyond the wave of enthusiasm generated by human-like responses that have been shown to pass the Turing Test [5] with a majority of human test subjects [6], the OpenAI’s GPT series demonstrated proficiency in academic fields such as physics in a number of benchmarks [7]. Both the initial version and later iterations, particularly GPT-4 [8], have achieved impressive results in physics, such as passing standardized exams, excelling in introductory courses and even coming close to passing entire degrees [9, 10, 11, 12, 13].
The technology is increasingly being embraced in physics education [14], offering promising applications for both teaching and learning physics. LLMs have proven valuable for teachers, helping in the creation of tailored materials and tasks, student assessments [15], and personalized feedback [16, 17, 18, 19, 20, 21], as well as for their training [22]. For students, these models represent ever-available, patient, and knowledgeable resources [23, 24]. However, these opportunities come with significant risks, particularly the potential for users to overly trust AI when assessing its scientific accuracy [25, 26].
Early publicly available generative AI systems have been limited to processing and generating text-based content only. Thus, earlier studies of AI’s performance on physics tasks, such as concept inventories, were restricted to text-based materials or textual descriptions of visual elements [9, 27, 28, 29, 30]. More recently, multimodal systems, which can also input and output auditory and visual data, have broadened the scope of such studies (e.g., [31]).
In this study, we confront GPT-4o [32], a popular AI model from OpenAI, with screenshots of tasks from physics concept inventories, which reflect the test items as seen by learners, including the accompanying diagrams, sketches, graphs, and illustrations. Earlier studies using GPT-4 and GPT-4o with the FCI [33], TUG-K [34] and BEMA [35] concept inventories demonstrated promising behavior, but also significant challenges, mainly related to the AI system’s limited ability to interpret visual information [30, 36, 37]. We considered the English version of each selected concept inventory and all its available translations.
@@ I.2 Concept inventories in physics
Examinations in physics generally involve symbolic and numerical calculations, whereas concept inventories are typically different, focusing primarily on conceptual understanding [38, 39]. Scores on these different types of assessments do not necessarily correlate, as for example, the FCI would under-predict success in a calculus-based physics course [40]. On the surface, this would work in AI’s favor, as AI systems used to be notoriously “bad at math,” which hampered their performance on traditional physics exam questions [9]; this is remedied in newer models which generate Python code for calculations or are explicitly “reasoning,” such as GPT-o1 and GPT-o3-mini [41, 42]. However, early investigations of AI’s performance on conceptual tasks suggest that these, too, can present several challenges [43, 10].
@@
Arguably, while many concept inventory items assess students’ understanding of core ideas and concepts [39], they do not necessarily capture evidence of scientific practices or crosscutting concepts [46] (even though this distinction is debatable [47]). It is thus important to emphasize that our study assesses AI’s performance on physics conceptual tasks, but does not evaluate if it has the qualities of a physicist.
@@ I.3 The language problem
Language plays a crucial role in learning physics. Expert physicists structure their knowledge using layered metaphorical systems and specific grammatical frameworks, which shape how they communicate complex ideas [48, 49]. These unconscious linguistic patterns can influence how students understand physics [50, 48]. Although English has become the lingua franca of most scientific communications, undergraduate students tend to be instructed in their native language, which is also the language in which they will tend to tackle physics problems.
@@
LLMs hold the promise of facilitating language-related learning tasks [51], yet they currently do not function equally well in all languages; for example, OpenAI’s research on the GPT-4’s performance shows inconsistent results across different languages [52]. Due to the disparities in the prevalence, quantity, and quality of information available across languages, there exists a disparity in the resources available for LLM training, which can have an impact on model performance. Nicholas and Bhatia highlight that although LLMs are designed to mitigate the issue of underrepresentation of certain languages in learning data, as of 2023, early LLMs were still predominantly trained on English materials [53]. While this may be changing as governments and companies attempt to strengthen AI capabilities in their countries (e.g., models like DeepSeek [54] or Qwen [55] developed by companies in China, or the Swiss AI Initiative [56]), it is likely that LLMs remain biased toward the needs of major economies with the financial resources needed to train AI systems. Similar issues are also discussed in the analysis conducted by “Cohere for AI” company in their report titled “The AI Language Gap” [57]. Feng et al. also acknowledge the disparity in LLM performance across different languages when it comes to abstaining from hallucinations, resulting in a gap of approximately 20% between high-resource and low-resource languages [58]; “hallucinations” refer to instances where an LLM generates plausible-sounding but inaccurate or entirely fabricated information, which are indicative of shortcomings in calibration and reasoning. Particularly relevant for our study is a recent finding that a Chinese-trained model performed better on the FCI when prompted in Chinese rather than English [29].
@@
Understanding and using physics-specific language is essential for physics literacy and learning. To address both the linguistic disparities of LLMs and the specialized nature of physics language, it is important to examine how LLMs perform across diverse languages in the context of physics education. While current research highlights the general challenges of multilingual use of LLMs, the intersection of language and subject-specific terminology, such as physics, is understudied and lacks a clear understanding of the current situation.
I.4 Relevance to Physics Education Research
While AI’s potential as a learning tool, assessment assistant, and research aid is widely acknowledged [59, 12, 60, 61, 62], the performance of LLMs on established and validated physics conceptual assessments, particularly in multilingual contexts, and on multiple subject domains, remains underexplored.
As AI assumes an increasingly large role in the educational process, it is also important for physics education stakeholders to develop a sense of its capabilities in physics-related tasks. Assessing AI’s ability to solve the kind of problems we use for assessing student understanding of physics concepts is a necessary step if we want to make appropriate and responsible use of these systems in physics education. The physics education research community, with its wealth of research-based assessment instruments, so-called concept inventories, is well positioned to engage in such evaluation. Concept inventories are developed to be robust tools for assessing university students’ conceptual understanding. Comparing AI’s performance to that of students can thus provide a student-centered reference point that is more meaningful for PER researchers than AI-facing benchmark assessments (which are oftentimes designed to assess physical reasoning tasks and inform machine-learning engineering [63, 64, 7, 65, 65, 66]), and they are more standardized and universal than exams from individual courses (e.g., [13]).
Comparing AI’s average performance to university students’ post-instruction performance can thus provide a rough student-centric measure of the level of performance of AI systems. There is a need for caution here, however. There are important differences in information processing in students and AI models. This means that, for example, when AI reaches a numerical performance similar to that of an average student, it does not necessarily mean that its strengths and difficulties are similar to those of an average student. In fact, previous studies have shown important differences in the profile of students’ and AI’s difficulties — i.e. AI’s displayed difficulties have been found to be uncharacteristic of typical student difficulties [37, 36].
One of the motivations for our work is exploring what types of tasks might be difficult for GPT-4o to solve. For instructors, knowing this would allow them to communicate to their students the strengths and weaknesses of the AI system, so that students may use it in responsible and productive ways when working unsupervised (e.g., while doing homework or project assignments) or in contexts beyond formal education. Exploring the different facets of an AI’s performance can also inform developers about AI’s potential as an educational tool for automated grading [67, 68, 20], generating feedback [69, 70], and personalizing instruction [71]. However, for AI to be genuinely useful in these applications, it must demonstrate consistency and reliability across diverse educational contexts. For example, a recent study suggests that an AI system’s ability to grade student answers on a topic is correlated to its problem-solving ability [68] on that topic. Analyzing GPT-4o’s performance on structured, standardized conceptual assessments, can therefore also provide insights into whether AI can reliably assist instructors in evaluating student understanding and be integrated into automated feedback and grading systems.
Another major motivation driving this research is the impacts of AI’s increasing accessibility on student learning and assessment validity in physics education [11]. With generative AI now readily available, students are already using it for study purposes, sometimes in ways that challenge traditional expectations of academic integrity [72]. More broadly, this issue ties into an ongoing discussion in PER regarding how assessments need to evolve in an AI-assisted learning environment and how physics curricula should adapt to the fact that many tasks can now be outsourced to AI.
Another key aspect of this study is its multilingual approach. While much of PER is conducted in English [73, 74], physics is taught in a wide range of languages, and the educational impact of AI should be considered beyond English-speaking classrooms. LMMs like GPT-4o have been trained primarily on English-language data [53], which raises concerns about whether their accuracy is consistent across different linguistic versions of physics assessments. If AI models perform significantly better in one language than in another, this could reinforce existing inequities in educational technology [75]. Understanding these potential disparities is necessary to ensure equitable access to AI-assisted learning tools and to assess whether AI can serve as a reliable resource for students who learn physics in languages other than English.
Lastly, our study contributes to a growing body of research on the multimodal performance of LLMs, particularly on tasks involving physics visual representations. This ties into the question of what types of tasks are especially difficult for AI to solve. While it is known that GPT-4o struggles with tasks involving kinematics graphs [31], it remains unclear if it experiences similar limitations across other types of physics problems and subject categories, and whether these weaknesses persist when the visual data are embedded in diverse real-world assessment formats.
This study is exploratory in nature; rather than aiming to provide definitive answers to broad questions about AI’s role in physics education, it seeks to identify emerging trends in AI performance on physics concept inventories. By analyzing GPT-4o’s accuracy on multiple-choice questions across different languages, with and without accompanying images, we aim to highlight patterns that warrant further investigation rather than making strong claims about AI’s conceptual reasoning abilities. Any broader generalizations about AI’s impact on assessment, instruction, or research would require a more detailed analysis of its reasoning processes and interactions with learners, which is beyond the scope of this paper. Nevertheless, by mapping out these initial trends, this study contributes to a growing conversation in PER about how AI is reshaping physics education, helping to inform future research directions and the responsible development and integration of AI into learning environments and educational processes.
I.5 Research questions
Despite a growing number of exploratory studies on LLMs in physics education [76, 15, 16, 17, 19, 20, 21, 30, 25, 26, 12, 77, 78, 79, 80, 81, 79], including some of our own [9, 37, 31], there remains a lack of a broad, systematic analysis of how multimodal AI models perform on validated, research-based physics assessments, particularly when these assessments include diverse subject domains, visual components, and multiple languages. Prior studies have often focused on individual inventories or English-only text-based inputs. This study extends that work by (i) synthesizing performance trends across inventories and physics subject categories, (ii) exploring performance in different languages, (iii) comparing AI performance to student benchmarks, and (iv) reevaluating AI’s visual interpretation challenges in a multilingual and multisubject framework. The following research questions guide our exploratory analysis:
@@ I.5 Research questions
RQ1. How does GPT-4o perform across different physics concept inventories in English?
RQ2. How does language influence the performance of the AI system?
RQ3. How does its performance compare to student performance at the undergraduate level?
RQ4. How does the presence of images influence the performance of the AI system?
These questions are, in turn, addressed in Section IV.
@@ II Data set
As our dataset, we used concept inventories published in PhysPort [82], containing a comprehensive collection of research-based inventories in multiple translations [83]. We included inventories that had at least a “bronze star” classification assigned on PhysPort. This indicates that they had been studied with respect to at least three of the seven research validation categories that the platform employs (research into student thinking; studied with student interviews, expert review, and appropriate statistical analysis; research conducted at multiple institutions, by multiple research groups, and with peer-reviewed publication). We excluded inventories for which we could not obtain an answer key. Figure 1 shows inventory items from the Force Concept Inventory in Persian and the Heat and Temperature Conceptual Evaluation in Chinese. These inventories cover a broad spectrum of physics subject categories, listed in Table 1.
Thirty-five languages are currently represented in PhysPort. While the authors are familiar with some of the languages, the quality of the majority of the non-English translations could not be evaluated.
Tables 2 through 5 show the concept inventories included in our investigation, along with their available translations and literature references. For some inventories, more than one version was available. The column “%Post” lists some post-instruction inventory scores reported in the literature for undergraduate-level courses; these are collected best-effort and not necessarily representative (typically and much more systematically, inventory gains are reported [84]). We omitted results with very small sample sizes, as well as reported scores at the graduate-student or post-doctoral level. For subsequent comparisons, the scores found for each inventory were averaged.
The rightmost column “Cat.” shows the subject-category identifier under which we classified the test. Explanations for the abbreviations used in this column are provided in Table 1. These labels are derived from the PhysPort classification, however, we divided electricity and magnetism into EM-F (primarily dealing with fields and potentials), and EM-C (primarily focusing on DC and AC circuits). The multi-subject Next Gen Physical Science Diagnostic (NGPSD) was given no classification.
@@
Chinese, English, Japanese, Portuguese, Spanish, Swedish | EM-F |
@@
Czech, English | MATH |
| CDPA | Concise Data Processing Assessment | Lab skills (uncertainty in measurement, relationship between functions graphs and numbers) | [91] | 2 | 39 [91] | English, Spanish | LAB |
@@
Chinese, English, Finnish, German, Greek, Spanish, Swedish | EM-C |
| DS | Density Survey | Mechanics content knowledge (density) | [98] | 1 | 57 [99] | English, German | MECH |
@@
English, Finnish, Indonesian, Swedish | MECH |
@@
Arabic, Bengali, Catalan, Chinese, Croatian, Czech, Dutch, English, Filipino, Finnish, French, German, Greek, Hebrew, Hindi, Hungarian, Icelandic, Italian, Japanese, Malay, Norwegian, Persian, Polish, Portuguese, Punjabi, Russian, Slovak, Spanish, Swedish, Tamil, Thai, Turkish | MECH |
| FMCE | Force and Motion Conceptual Evaluation | Mechanics content knowledge (kinematics, forces, energy, graphing) | [109] | v99 | 55 [110] | English, Indonesian, Japanese, Spanish | MECH |
| FORT | Montana State University Formal Reasoning Test | Scientific reasoning (hypothesis testing, correlational reasoning, probability, control of variables, proportional reasoning) | [111] | 1 | 55 [111] | English | REAS |
| FTGOT | Four-tier Geometrical Optics Test | Waves / Optics content knowledge (plane mirrors, spherical mirrors, lenses); used as two-tier test | [112] | 1 | 18 [112] | English, Turkish | OPT |
@@
English | MECH |
| GECI | Greenhouse Effect Concept Inventory | Astronomy content knowledge (types of greenhouse gases, types of electromagnetic energy, energy equilibrium balance, greenhouse effect mechanisms, global warming vs. greenhouse effect) | [114] | vC | 55 [114] | English, Japanese | AST |
| HTCE | Heat and Temperature Conceptual Evaluation | Thermal / Statistical content knowledge (temperature, phase change, heat transfer, thermal properties of materials) | [115] | 1 | 78 [115] | Chinese, English | THERM |
@@
English: 26
French: 31 [116] |
English, French | MECH |
| LPCA | Light Phenomena Conceptual Assessment | Waves / Optics content knowledge (reflection, refraction, Snells law, wavelength and frequency, light scattering, electromagnetic spectrum, the human eye) | [117] | 1 | 41 [117] | English | OPT |
@@
English, Finnish, French, German, Greek, Italian, Japanese, Malay, Persian, Portuguese, Spanish, Turkish | MECH |
@@
English | EM-F |
| MUQ | Measurement Uncertainty Quiz | Lab skills (calculating error from measurements, accuracy and precision, sources of error) | [126] | 1 | — | English | LAB |
| MWCS | Mechanical Wave Conceptual Survey | Waves / Optics content knowledge (mechanical waves, wave propagation, wave superposition, wave reflection, standing waves) | [127] | 1 | 47 [128] | English, Spanish, Thai | OPT |
| 2 | English, Spanish | ||||||
@@
Arabic, English | AST |
| NGPSD | Next Gen Physical Science Diagnostic | Mechanics content knowledge (magnetism, static electricity, energy, forces, waves and sound, light) | [130] | 2 | — | English | — |
@@
| PIQL | Physics Inventory of Quantitative Literacy | Mathematics scientific reasoning (proportional reasoning, reasoning with signed quantities, co-variational reasoning) | [131] | 2.2 | 55 [131] | English | REAS |
| QMCA | Quantum Mechanics Concept Assessment | Modern / Quantum Content knowledge (wave functions, measurement, time dependence, probability, infinite square well, 1D tunneling, energy levels, spins) | [132] | 5.5.7 | 54 [132] | English, Portuguese | QP |
| 6.6.2 | English | ||||||
@@
English, Spanish | QP |
| QMS | Quantum Mechanics Survey | Modern / Quantum content knowledge (wave functions, measurement, expectation values, Hamiltonian, time dependence, probability, infinite square well, finite square well, harmonic oscillator, 1D tunneling) | [136] | 18 | 38 [136] | English | QP |
@@
English | EM-F |
| RCI | Relativity Concept Inventory | Modern / Quantum content knowledge | [140] | 1 | 71 [140] | English | RELA |
@@
English, French | MECH |
| RRMCS | Rotational and Rolling Motion Conceptual Survey | Mechanics content knowledge (rotational kinetic energy, torque, rotational kinematics, moment of inertia) | [144] | 1 | 75 [144] | English | MECH |
| SGCE | Symmetry and Gauss’s Law Conceptual Evaluation | Electricity / Magnetism content knowledge (symmetry, electric field, electric flux) | [145] | 1 | 49 [145] | English | EM-F |
@@
English, Japanese, Spanish | AST |
| STPFaSL | Survey of Thermodynamic Processes and First and Second Laws | Thermal / Statistical content knowledge (first law of thermodynamics, second law of thermodynamics, PV diagrams, reversible processes, irreversible processes) | [147] | short | 37 [147] | Chinese, English, Indonesian, Portuguese | THERM |
| long | English | ||||||
| TCE | Thermal Concept Evaluation | Thermal / Statistical content knowledge (temperature, heat transfer, phase change, thermal properties of materials) | [148] | 1 | 78 [148] | Chinese, English, Japanese, Portuguese | THERM |